# Neuro-inspired Ensemble-to-Ensemble Communication Primitives for Sparse and Efficient ANNs

### v1

@@ Abstract
Index Terms:
Neuronal Ensembles, Sparsity in Deep Learning, Functional Connectivity, Computer Vision
@@ II Background: Sparsity in Deep Learning
Our work builds on the growing body of research of brain-inspired deep learning and sparsity in ANNs by incorporating functional connectivity patterns from biological networks as a structural bias in ANN architectures. Our sparse architecture builds on neuronal ensembles, i.e., groups (used interchangeably in the paper) within each layer and their inter-layer communication. This introduces the notion of pathways: active groups that transmit information between layers, bringing modularity and structured sparsity into the design. In addition, we explore the integration of dynamic sparse training techniques within this biologically informed framework. We introduce a Hebbian-based criterion for pruning and growing connections inspired by biological synaptic plasticity.
@@ III Methodology
We will examine the role of the ensemble-to-ensemble information flow observed in biological networks by imposing a structural bias on the feedforward layers of ANNs. A key challenge is how the functional grouping observed in biological neurons can be translated into structural grouping in ANNs. We propose that grouping can be achieved by ensuring that neurons within the same group share highly similar connectivity patterns, that is neurons in a specific group (say i) connect to the same neurons in the next layer with the same probability. Additionally, each group is strongly coupled with its corresponding group in the next layer, while maintaining only sparse connections with other groups in that layer. This design results in a highly sparse network with small-world-like properties across its layers. In the following paragraphs, we present the details of our architecture.
@@ III-A Sparse layer structure inspired by primary visual cortex (V1)
    For each neuron in group $k$ of layer $i$, connect it to neurons in group $k$ of layer $i+1$ with probability p, and to neurons in all other groups of layer $i+1$ with a much lower probability $\textbf{p'}\ll\textbf{p}$.
@@ III-B Grouping strategies
Several grouping strategies can be considered: neurons can be grouped by their index (”Index-based”, Fig. 3A), preserving the ordering induced by the input. Alternatively, neurons can be assigned to groups randomly, disregarding any structure (”Random”, Fig. 3B). In this work, we propose a hybrid method inspired by the MLPMixer architecture [62], which we call the Mixer strategy (”Mixer” Fig. 3C). The Mixer alternates between two grouping schemes in successive layers:
@@ III-D Rewiring Criteria during Training
TABLE I: Test accuracies (%) of different connectivity patterns on FashionMNIST (F-MNIST), CIFAR-10 and CIFAR-100. We compare fully-connected architectures with the same number of parameters (v1) or same width (v2) as ours, as well as ER random sparse networks and our V1-inspired topologies.
Connectivity Pattern F-MNIST Accuracy (%) CIFAR-10 Accuracy (%) CIFAR-100 Accuracy (%) Parameters
@@
G2GNet (Index Grouping, Ours) 90.4 70.9 42 806,100
G2GNet (Mixer Grouping, Ours) 91.14  71.5 42.2 806,100
@@ IV-A Architectural Details
Figure 4: Experimental results on CIFAR-10: (a) As the inter-pathway probability $p^{\prime}$ of connecting sparsely connected groups, and thus the number of edges, increases, accuracy improves and plateaus around 70.8% for $p^{\prime}\in[0.08,0.15]$, corresponding to approximately 18% to 25% sparsity. Compared to Erdős-Rényi (ER) graphs of equal number of edges, our structure bias consistently yields higher accuracy. (b) The Mixer technique outperforms grouping by indexing across a wide range of $p^{\prime}$ values, while random grouping significantly underperforms. (c) Increasing the intra-pathway probability $p$ of connecting densely connected groups leads to higher accuracy; we therefore set $p=1$ in all experiments. (d) Effect of the rewiring update interval: accuracy is evaluated as a function of the number of training iterations between edge removal and addition steps. The best performance is observed when the update occurs every 1000 iterations. (e) Effect of the $\%$ of drop/add edges: accuracy is measured for different proportions of weights pruned and regrown, with the update interval fixed at 1000 iterations. The optimal performance is achieved around a 2.5% update rate. Reported values are mean accuracies over five runs with identical hyperparameters. For grouping strategy comparisons, the connectivity matrix initialization is kept fixed.
@@
Comparing grouping techniques In Figure 4B, we also report the impact of the different grouping techniques described in Section III-B. We find that the Mixer technique performs best, as it enables dense modules to process information from diverse regions of the input image. Interestingly, when grouping is random, the information flow lacks any structure related to the spatial proximity of the input, and the results deteriorate to levels like the Erdős–Rényi connectivity pattern.
@@ IV-B Sparse Training Evaluation
Table II show the performance of various dynamic sparse training (DST) strategies applied to G2GNet on the CIFAR-10 and CIFAR-100 datasets. We evaluate all combinations of pruning (edge removal) criteria and the two growth (edge addition) methods described in Section III. Specifically, edges are pruned using one of the following strategies: (i) random selection, (ii) magnitude-based pruning (following the SET criterion [52]), or (iii) our proposed Hebbian-based criterion. Edges are then regrown either randomly or using the Hebbian criterion, which prioritizes connections between neurons with highly correlated activation patterns.
We do not consider DST methods that rely on dense gradient storage [55, 64], as they require maintaining gradient information for inactive connections, which undermines the computational benefits of sparsity.
Empirically, we find that DST strategies consistently improve the performance of G2GNet on CIFAR-10, and on CIFAR-100 when the Hebbian growth criterion is employed. Among the various combinations, the Hebbian criterion generally performs best for edge addition, consistently outperforming random growth. The most effective configuration involves pruning by magnitude (SET) and growing with the Hebbian criterion on CIFAR-10, while on CIFAR-100, the best results are achieved by using the Hebbian criterion for both pruning and growth.
We then fix the DST configuration to use Hebbian-based criteria for both pruning and growth, and conduct a sensitivity analysis on two key hyperparameters: the weight update interval and the proportion of connections modified per update (see Figures 4D, 4E). In line with prior work [63], we find that performance is optimized when updates occur every 1000 training iterations, with an update rate of approximately 2.5%. When the update interval is too short or the update rate too high, the network struggles to adapt and train the newly added parameters effectively. On the other hand, if updates are too infrequent or the update rate too low, the impact of the rewiring mechanism diminishes.
@@ V Discussion and Future Work
The initial connectivity pattern imposed by G2GNet may not contain the “winning ticket”, a subnetwork that performs as well as the full network, as described by the Lottery Ticket Hypothesis. To address this, we incorporate a dynamic edge rewiring mechanism that maintains a constant parameter count while adapting connectivity during training. Central to this mechanism is our proposed Hebbian-inspired criterion, which adds or prunes edges based on the activation correlation between connected neurons. Our approach yields strong performance on computer vision benchmarks and enables structured sparsity, a form of sparsity that is more compatible with current hardware acceleration.
@@
While our method is only conceptually inspired by biological principles, such as ensemble-to-ensemble communication and information flow, it opens promising directions for future research. In particular, the notion of functional groups in biological networks could be extended to dynamic group formation in ANNs, driven by input data, internal network state, or even as a means for network training and avoiding of catastrophic forgetting. Currently, our experiments focus on feedforward layers, which limits the complexity of tasks evaluated. However, the underlying structural bias of G2GNet is applicable to convolutional layers as well as to the query, key, and value matrices of attention mechanisms in transformer-based architectures. G2GNet also provides an example of how the organizational principles identified in neuroscience (e.g., in [42]) may inform the design of ANNs, providing also a simulation tool for testing of neuro-computational hypotheses.

### v2

@@ II Background: Sparsity in Deep Learning
Our work builds on the growing body of research of brain-inspired deep learning and sparsity in ANNs by incorporating functional connectivity patterns from biological networks as a structural bias in ANN architectures. Our sparse architecture builds on neuronal ensembles, i.e., groups (used interchangeably in the paper) within each layer and their inter-layer communication. This introduces the notion of pathways: active groups that transmit information between layers, bringing modularity and structured sparsity into the design. In addition, we explore the integration of dynamic sparse training techniques within this biologically informed framework. We introduce a Hebbian-like criterion for pruning and growing connections inspired by biological synaptic plasticity.
@@ III Methodology
We will examine the role of the ensemble-to-ensemble information flow observed in biological networks by imposing a structural bias on the feedforward layers of ANNs. A key challenge is how the functional grouping observed in biological neurons can be translated into structural grouping in ANNs. We propose that grouping can be achieved by ensuring that neurons within the same group share highly similar connectivity patterns, that is, neurons in a specific group (say i) connect to the same neurons in the next layer with the same probability. Additionally, each group is strongly coupled with its corresponding group in the next layer, while maintaining only sparse connections with other groups in that layer. This design results in a highly sparse network with small-world-like properties across its layers. In the following paragraphs, we present the details of our architecture.
@@ III-A Sparse layer structure inspired by primary visual cortex (V1)
    For each neuron in group $k$ of layer $i$, connect it to neurons in group $k$ of layer $i+1$ with probability p, and to neurons in all other groups of layer $i+1$ with a much lower probability $\textbf{p}^{\prime}\ll\textbf{p}$.
@@ III-B Grouping strategies
Several grouping strategies can be considered: neurons can be grouped by their index (“Index-based”, Fig. 3A), preserving the ordering induced by the input. Alternatively, neurons can be assigned to groups randomly, disregarding any structure (“Random”, Fig. 3B). In this work, we propose a hybrid method inspired by the MLPMixer architecture [62], which we call the Mixer strategy (“Mixer” Fig. 3C). The Mixer alternates between two grouping schemes in successive layers:
@@ III-D Rewiring Criteria during Training
TABLE I: Test accuracies (%) of different connectivity patterns on FashionMNIST (F-MNIST), CIFAR-10 and CIFAR-100. We compare fully-connected architectures with the same number of parameters (v1) or same width (v2) as ours, as well as ER random sparse networks and our V1-inspired topologies. e
Connectivity Pattern F-MNIST Accuracy (%) CIFAR-10 Accuracy (%) CIFAR-100 Accuracy (%) Num. Parameters
@@
G2GNet (Proposed) 91.14 71.5 42.2 806,100
@@ IV-A Architectural Details
Figure 4: Experimental results on CIFAR-10 (a) As the inter-pathway probability $p^{\prime}$ of connecting sparsely connected groups, and thus the number of edges, increases, accuracy improves and plateaus around 70.8% for $p^{\prime}\in[0.08,0.15]$, corresponding to approximately 18% to 25% sparsity. Compared to Erdős-Rényi (ER) graphs of equal number of edges, our structure bias consistently yields higher accuracy. (b) The Mixer technique outperforms grouping by indexing across a wide range of $p^{\prime}$ values, while random grouping significantly underperforms. (c) Increasing the intra-pathway probability $p$ of connecting densely connected groups leads to higher accuracy; we therefore set $p=1$ in all experiments. (d) Effect of the rewiring update interval: accuracy is evaluated as a function of the number of training iterations between edge removal and addition steps. The best performance is observed when the update occurs every 1000 iterations. (e) Effect of the $\%$ of drop/add edges: accuracy is measured for different proportions of weights pruned and regrown, with the update interval fixed at 1000 iterations. The optimal performance is achieved around a 2.5% update rate. Reported values are mean accuracies over five runs with identical hyperparameters. For grouping strategy comparisons, the connectivity matrix initialization is kept fixed.
@@
Comparing grouping techniques In Figure 4B, we also report the impact of the different grouping techniques described in Section III-B. We found that the Mixer technique performs best, as it enables dense modules to process information from diverse regions of the input image. Interestingly, when grouping is random, the information flow lacks any structure related to the spatial proximity of the input, and the results deteriorate to levels like the Erdős–Rényi connectivity pattern.
@@ IV-B Sparse Training Evaluation
Table II shows the performance of various dynamic sparse training (DST) strategies applied to G2GNet on the CIFAR-10 and CIFAR-100 datasets. We evaluated all combinations of pruning (edge removal) criteria and the two growth (edge addition) methods described in Section III. Specifically, edges are pruned using one of the following strategies: (i) random selection, (ii) magnitude-based pruning (following the SET criterion [52]), or (iii) our proposed Hebbian-based criterion. Edges are then regrown either randomly or using the Hebbian criterion, which prioritizes connections between neurons with highly correlated activation patterns.
We did not consider DST methods that rely on dense gradient storage [55, 56], as they require maintaining gradient information for inactive connections, which undermines the computational benefits of sparsity.
Empirically, we found that DST strategies consistently improve the performance of G2GNet on CIFAR-10, and on CIFAR-100 when the Hebbian growth criterion is employed. Among the various combinations, the Hebbian criterion generally performs best for edge addition, consistently outperforming random growth. The most effective configuration involves pruning by magnitude (SET) and growing with the Hebbian criterion on CIFAR-10, while on CIFAR-100, the best results are achieved by using the Hebbian criterion for both pruning and growth.
We then fixed the DST configuration to use Hebbian-based criteria for both pruning and growth, and conducted a sensitivity analysis on two key hyperparameters: the weight update interval and the proportion of connections modified per update (see Figures 4D, 4E). In line with prior work [63], we found that performance is optimized when updates occur every 1000 training iterations, with an update rate of approximately 2.5%. When the update interval is too short or the update rate too high, the network struggles to adapt and train the newly-added parameters effectively. On the other hand, if updates are too infrequent or the update rate too low, the impact of the rewiring mechanism diminishes.
@@ V Discussion and Future Work
The initial connectivity pattern imposed by G2GNet may not contain the “winning ticket”, a subnetwork that performs as well as the full network, as described by the Lottery Ticket Hypothesis. To address this, we incorporate a dynamic edge rewiring mechanism that maintains a fixed number of pruned and newly-activated edges while adapting connectivity during training. Central to this mechanism is our proposed Hebbian-inspired criterion, which adds or prunes edges based on the activation correlation between connected neurons. Our approach yields strong performance on computer vision benchmarks and enables structured sparsity, a form of sparsity that is more compatible with current hardware acceleration.
@@
While our method is only conceptually inspired by biological principles, such as ensemble-to-ensemble communication and information flow, it opens promising directions for future research. In particular, the notion of functional groups in biological networks could be extended to dynamic group formation in ANNs, driven by input data, internal network state, or even as a means for network training and avoiding of catastrophic forgetting. Currently, our experiments focus on feedforward layers, which limits the complexity of tasks evaluated. However, the underlying structural bias of G2GNet is applicable to convolutional layers as well as to the query, key, and value matrices of attention mechanisms in transformer-based architectures. Moreover, part of our on-going efforts includes the analysis of these ideas on SNNs and an in-depth examination of mixer-like grouping strategies. G2GNet provides an example of how the organizational principles identified in neuroscience (e.g., in [42]) may inform the design of ANNs, providing also a simulation tool for testing of neuro-computational hypotheses.
Acknowledgments
This work has been partially supported by the project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. It has also received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 101007926 neuronsXnets as well as from the Hellenic Foundation Research Institute (HFRI) with the neuron-AD project number 04058 and project number 2285 (PI: Maria Papadopouli).